# VideoGen-of-Thought: A Collaborative Framework for Multi-Shot Video Generation

### v1

@@ Abstract
Current video generation models excel at generating short clips but still struggle with creating multi-shot, movie-like videos. Existing models trained on large-scale data on the back of rich computational resources are unsurprisingly inadequate for maintaining a logical storyline and visual consistency across multiple shots of a cohesive script since they are often trained with a single-shot objective. To this end, we propose VideoGen-of-Thought (VGoT), a collaborative and training-free architecture designed specifically for multi-shot video generation. VGoT is designed with three goals in mind as follows. Multi-Shot Video Generation: We divide the video generation process into a structured, modular sequence, including (1) Script Generation, which translates a curt story into detailed prompts for each shot; (2) Keyframe Generation, responsible for creating visually consistent keyframes faithful to character portrayals; and (3) Shot-Level Video Generation, which transforms information from scripts and keyframes into shots; (4) Smoothing Mechanism that ensures a consistent multi-shot output. Reasonable Narrative Design: Inspired by cinematic scriptwriting, our prompt generation approach spans five key domains, ensuring logical consistency, character development, and narrative flow across the entire video. Cross-Shot Consistency: We ensure temporal and identity consistency by leveraging identity-preserving (IP) embeddings across shots, which are automatically created from the narrative. Additionally, we incorporate a cross-shot smoothing mechanism, which integrates a reset boundary that effectively combines latent features from adjacent shots, resulting in smooth transitions and maintaining visual coherence throughout the video. Our experiments demonstrate that VGoT surpasses existing video generation methods in producing high-quality, coherent, multi-shot videos. Project page: https://cheliosoops.github.io/VGoT
Figure 1: Illustration of VideoGen-of-Thought (VGoT). (a) Comparison of existing methods with VGoT in multi-shot video generation. Existing methods struggle with maintaining consistency and logical coherence across multiple shots, while VGoT effectively addresses these challenges through a multi-shot generation approach. (b) Overview of our proposed framework VGoT, which is consist of the Script Module which generates detailed shot descriptions from five domains, the KeyFrame Module to create keyframes from scripts, the Shot-Level Video Module which synthesizes video latents on conditional with keyframes and scripts, and the Smooth Module ensures seamless transitions across shots, resulting in a cohesive video narrative.
@@ 1 Introduction
Recent advancements in video generation techniques have yielded impressive results, particularly in creating short, visually appealing clips [6, 8, 9, 69, 28]. These advancements have been powered by increasingly sophisticated generative models, ranging from diffusion models [29, 56, 49, 6] to auto-regressive models [20, 65, 39, 62], supported by large-scale datasets [31, 50, 51]. These methods have enabled the generation of high-quality, realistic short videos. However, generating multi-shot videos from a brief user input script remains a substantial challenge. Despite existing works successfully achieving long-term video generation, the primary focus has been on extending the duration of single-shot videos [19, 40, 61, 35]. In contrast, multi-shot video generation involves additional complexities in ensuring coherence across different shots. Moreover, while simply scaling up models or using larger datasets, as done in single-shot long video generation methods [63, 45], can enhance the visual quality of individual clips, it requires vast computational resources and extensive training times. More importantly, they fail to solve the challenges inherent in multi-shot video generation, which is a more complex task that requires maintaining reasonable logic in the narrative, and visual consistency across different scenes—all of which are essential for producing multi-shot, movie-like content.
The limitations of existing approaches underscore a fundamental gap in the field: increasing model size and dataset volume alone is insufficient for generating reasonable and consistent videos with multiple shots. The primary challenge is not solely to achieve high visual fidelity but also to ensure that each shot seamlessly connects to the next, thereby contributing to a coherent and compelling narrative structure. When applied to multi-shot video generation, we observed that current methods often fall short in maintaining a consistent storyline and visual dynamics across multiple shots [8, 9, 69]. The absence of a structured approach for organizing and managing the generation process often results in a breakdown of narrative flow, with logical transitions between scenes being lost. Therefore, there is a pressing need for an approach that effectively balances visual quality with narrative consistency, managing the generation process without relying purely on computational scaling.
Inspired by the “thought process” behind the creation of real movies, we propose VideoGen-of-Thought (VGoT)—a collaborative framework designed to tackle these challenges and create multi-shot videos in a step-by-step manner. An existing approach with a similar manner called “Chain of Thought” [64] has been proven efficient in Large Language Models (LLMs) [1, 15] and some visual perception tasks [18]. In this paper, we extend this approach to video generation by breaking down the complex process into four distinct yet collaborative modules as shown in Fig. 1: Script Generation Module converts an simple and high-level story description into detailed prompts for each shot to provide a solid narrative foundation; Keyframe Generation Module generates visually consistent keyframes based on the script, ensuring consistency in the appearance of characters, environments, and objects; Shot-Level Video Module produces video latents from corresponding keyframes and prompts at each shot that represents a shot-level video clip; finally, the Smooth Module integrates a cross-shot smoothing mechanism that processes information embeddings for the upcoming shot when the current one is nearing completion, ensuring a seamless connection facilitated by a reset boundary. This mechanism provides a smooth transition between shots, resulting in an end-to-end shot-by-shot video generation process that maintains visual and narrative continuity across the entire video.
As a text-to-video [53, 57] (T2V) architecture, our method VGoT employs a structured prompt design approach inspired by cinematic scriptwriting to ensure the generated videos are reasonable. Given a one-sentence script as user input, our method first prepares it into short descriptions for each shot and then extends them into five domains including character, background, relations, camera pose, and High Dynamic Range (HDR) descriptions—which together form a comprehensive movie-like script, as shown in the left part of Fig 2. Reasonability here refers to the logical flow and narrative coherence of the generated video, ensuring that each shot contributes meaningfully to the storyline. Furthermore, our method also tackles the issue of consistency across shots by introducing mechanisms for both temporal and visual coherence. Specifically, we use identity-preserving (IP) embeddings from avatars automatically derived from the narrative to maintain consistent character portrayal throughout the video. Additionally, our cross-shot smoothing mechanism, which leverages latent features from adjacent shots, ensures seamless transitions, resulting in a cohesive and consistent multi-shot video.
Our experimental results demonstrate the efficacy of VGoT in generating high-quality, logically consistent, multi-shot videos. We evaluated VGoT using ten different stories, and a fair comparison with existing baselines highlights the superior performance of our approach in terms of narrative coherence, visual quality, and cross-shot consistency. The main contributions of this paper are as follows:
@@
    We propose a collaborative, training-free architecture VideoGen-of-Thought for multi-shot video generation, addressing the key challenges of reasonability and consistency across multiple shots.
@@
    We introduce a structured prompt design that emulates cinematic scriptwriting, ensuring logical coherence and narrative depth in generated videos.
@@
    We develop a cross-shot consistency mechanism using identity-preserving embeddings and a cross-shot smoothing technique to maintain visual continuity and temporal coherence across different shots.
@@ 2 Related Work
Video generation has progressed significantly following the great success of diffusion models, leading to two important research categories: long video synthesis and multi-shot story generation. These areas focus on generating high-quality, consistent videos either as extended single shots or as coherent sequences across multiple scenes.
Long Video Synthesis. Long video synthesis has advanced through diffusion-based methods and autoregressive approaches. Diffusion models based on the Latent Diffusion Model [49] utilize iterative refinement to generate visually consistent frames and have been effective for short sequences and other domains [26, 6, 8, 9, 68, 66, 7, 25, 71, 79, 22, 32, 44, 16, 76, 54, 78, 82, 46, 41, 42, 58, 80, 43, 38]. However, they struggle with maintaining coherence over extended video lengths. Autoregressive methods [20, 36, 73] predict frames sequentially but often face error accumulation, making long-term consistency challenging and computationally expensive. Additionally, training-free methods like FIFO-Diffusion [33] generate long sequences without training but lack mechanisms to manage transitions across shots, limiting their effectiveness in narrative-driven content. Overall, while these approaches achieve visual fidelity, they fail to ensure logical coherence across extended sequences. In contrast, VideoGen-of-Thought (VGOT) leverages a modular approach that includes cross-shot smoothing mechanisms to ensure both visual consistency and narrative coherence, offering a more holistic solution for generating long-form videos.
Multiple Shot Story Generation. Multiple Shot Story Generation focuses on maintaining narrative coherency across distinct scenes. Animate-a-Story [27] uses retrieval-augmented generation to ensure visual consistency but struggles with maintaining logical narrative transitions. DreamFactory [67] employs a multi-agent system for collaboration across shots, but lacks robust mechanisms for logical continuity, leading to inconsistencies. MovieDreamer [81] uses a hierarchical framework to manage coherence but relies heavily on predefined structures, limiting adaptability to dynamic stories. Flexifilm [40] attempts to address scene adaptability by incorporating flexible conditions, but it still falls short in ensuring logical cohesion between shots and often requires manual adjustments for consistency. Collectively, these methods often require significant manual intervention to achieve a consistent narrative, limiting scalability for fully automated storytelling. In contrast, our work provides a detailed and integrated framework that not only ensures cross-shot coherence through a novel smoothing mechanism but also employs rigorous evaluation methods to achieve both logical consistency and scalability, setting a new paradigm for multi-shot video generation.
@@ 3 Preliminaries and Problem Formulation
Figure 2: The FlowChart of VideoGen-of-Thought. Left: Shot descriptions are generated based on user prompts, describing various attributes including character details, background, relations, camera pose, and lighting HDR. Pre-shot descriptions provide a broader context for the upcoming scenes. Middle Top: Keyframes are generated using a text-to-image diffusion model conditioned with identity-preserving (IP) embeddings, which ensures consistent representation of characters throughout the shots. IP portraits help maintain visual identity consistency. Right: The shot-level video clips are generated from keyframes, followed by shot-by-shot smooth inference to ensure temporal consistency across different shots. This collaborative framework ultimately produces a cohesive narrative-driven video.
@@ 3.1 Preliminaries
Diffusion Models [29, 55, 56] are generative models that trained to approximate data distribution $p(x)$ from random noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. In the forward process, noise scheduled by the variance $\beta_{t}$ is gradually added as $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})$ over $T$ time steps, resulting in noisy latent variables $x_{T}$. Then a learnable model $\mu_{\theta}$ parameterized by $\theta$ aims to denoise $x_{T}$ back to the original data sample $x_{0}$, with a parameterized distribution given by $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$ where $\mu_{\theta}$ and $\Sigma_{\theta}$ are predicted mean and variance. The objective is to minimize the difference between predicted and true noise through the loss function:
$\mathcal{L}_{\text{uncond diff}}=\mathbb{E}_{x_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I}),t}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right]$ (1)
In the reverse process, an overall sample queue is maintained as $[x_{T};x_{T-1};\dots;x_{0}]$ at the pixel level. Latent diffusion models [49] successfully map this process into a lower-dimensional latent space using a variational autoencoder (VAE) [34, 70, 10, 37]. This transformation significantly enhances flexibility, allowing additional conditions to be incorporated through cross-attention mechanisms [60], which can be formulated as:
@@
where $Q$ is derived from the noisy image embeddings $e_{i}$, and $K$ and $V$ are derived from the text embeddings $e_{t}$ in text-to-image tasks. This mechanism has been further extended to the temporal dimension in conditional diffusion models for video generation [26, 6, 8, 74, 28, 75, 77]. For a video $V\in\mathbb{R}^{F\times H\times W\times 3}$ with $F$ frames and spatial resolution $H\times W$, a series of latent codes $\{z^{f}\}_{f=0}^{F-1}$, where $z^{f}\in\mathbb{R}^{h\times w\times d}$ with $d$ channels will be denoise in the same level. These video diffusion models, denoted as $\epsilon_{\theta}(z_{t},t,c)$, are capable of predicting $\hat{z}$, which can be decoded into video frames based on given conditions $c$.
While most existing methods aim to achieve long video generation through enlarging the shape of latent code $z$ and the size of model parameters $\theta$, FIFO-Diffusion [33] provides a new approach through denoising a queue of latents $Q_{k}=\{z^{f}\}_{f_{k}}^{f_{k}+n}$, updating $Q$ with DDIM[55] sampler $\Phi$ as:
$Q_{k}\leftarrow\Phi(Q_{k},\tau_{k},c;\epsilon_{\theta})$ (3)
for each timestep $\tau_{k}\in[0,T-n]$ from different level, in effect implementing long-term frame-by-frame prediction. Despite FIFO excelling in frame-by-frame prediction, it fails to handle the quick transition for different shots in multi-shot video generation. Moreover, as a text guided inference method, FIFO can’t receive conditions like image. In our work, we use the generative capability of diffusion models and process cross-shot smoothness with a FIFO-like mechanism at the shot level.

### v2

# VideoGen-of-Thought: Step-by-step generating multi-shot video with minimal manual intervention

@@ Abstract
Current video generation models excel at short clips but fail to produce cohesive multi-shot narratives due to disjointed visual dynamics and fractured storylines. Existing solutions either rely on extensive manual scripting/editing or prioritize single-shot fidelity over cross-scene continuity, limiting their practicality for movie-like content. We introduce VideoGen-of-Thought (VGoT), a step-by-step framework that automates multi-shot video synthesis from a single sentence by systematically addressing three core challenges: (1) Narrative Fragmentation: Existing methods lack structured storytelling. We propose dynamic storyline modeling, which first converts the user prompt into concise shot descriptions, then elaborates them into detailed, cinematic specifications across five domains (character dynamics, background continuity, relationship evolution, camera movements, HDR lighting), ensuring logical narrative progression with self-validation. (2) Visual Inconsistency: Existing approaches struggle with maintaining visual consistency across shots. Our identity-aware cross-shot propagation generates identity-preserving portrait (IPP) tokens that maintain character fidelity while allowing trait variations (expressions, aging) dictated by the storyline. (3) Transition Artifacts: Abrupt shot changes disrupt immersion. Our adjacent latent transition mechanisms implement boundary-aware reset strategies that process adjacent shots’ features at transition points, enabling seamless visual flow while preserving narrative continuity. By integrating these innovations into a training-free pipeline, VGoT generates multi-shot videos that outperform state-of-the-art baselines by 20.4% in within-shot face consistency and 17.4% in style consistency, while achieving over 100% better cross-shot consistency and 10× fewer manual adjustments than alternatives like MovieDreamer and DreamFactory. Our work redefines automated long-video generation, bridging the gap between raw visual synthesis and director-level storytelling.
Figure 1: Illustration of VideoGen-of-Thought (VGoT). (a) Comparison of existing methods with VGoT in multi-shot video generation. Existing methods struggle with maintaining reasonability and consistency across multiple shots, while VGoT effectively addresses these challenges through a multi-shot generation approach. (b) Challenges solved by VGoT: addressing narrative fragmentation with dynamic storylines modeling across five domains (characters/backgrounds/relations/camera/HDR), tackling visual inconsistency via identity-aware cross-shot propagation to create keyframes using IPP tokens derived from narrative elements, and solving transition artifacts during multi-shot video synthesizes through adjacent latent transition mechanisms.
@@ 1 Introduction
Recent advancements in video generation techniques have yielded impressive results, particularly in creating short, visually appealing clips [5, 7, 8, 53, 24]. These advancements have been powered by increasingly sophisticated generative models, ranging from diffusion models [25, 46, 40, 5] to auto-regressive models [15, 49, 33, 48], supported by large-scale datasets [28, 41, 42]. These methods have enabled the generation of high-quality and realistic short videos. However, generating multi-shot videos from a brief user input script remains a substantial challenge. Unlike single-shot video generation, which focuses on creating a coherent clip from a single prompt, multi-shot video generation requires the model to maintain both reasonable storylines and visual consistency across multiple shots. This task involves additional complexities, such as ensuring logical transitions between scenes and maintaining consistent appearance (e.g., character identity, overall style, etc.) throughout the video. Current video generation methods [7, 8, 53, 27, 54] often fall short in these areas, resulting in fragmented narratives and inconsistent visual elements across shots. Furthermore, sometimes real movies require the same character to appear in different ways based on the storylines, which should be faithful to the same identity but not the same traits (e.g., expression, appearance, relationships, etc.). The requirement for high-level identity preservation [55, 57, 62] across shots remains an open question, which is essential for cross-shot consistency.
Existing multi-shot video generation approaches suffer from several limitations. For instance, methods like MovieDreamer [60] require plenty of manual input, including mountains of script writing (e.g., character appearance, scene elements, detailed plots, etc.) and image selection. Other approaches, such as DreamFactory [51], focus on multi-agent pipelines but require specific documents and repeated manual adjustment for each story, restricting the capability of easy usage. The need for heavy manual intervention not only increases the workload but also limits their practicality for automated movie-like content creation. In contrast, our work aims to decompose the complex task of multi-shot video generation into smaller, manageable problems and solve them in a step-by-step manner with minimal manual intervention, as shown in Figure 1. Our approach proposes to systematically and automatically address three core challenges: (1) narrative fragmentation through dynamic storyline modeling; (2) visual inconsistency via identity-aware cross-shot propagation; and (3) transition artifacts using adjacent latent transition mechanisms.
We propose VideoGen-of-Thought (VGoT), an end-to-end framework that generates multi-shot video with reasonable storylines and visual consistency from one sentence with minimal manual intervention. Our framework addresses three fundamental challenges through a structured pipeline as shown in Fig 2: First, VGoT tackles narrative fragmentation through converting a brief user prompt into short descriptions for across shots to obtain a reasonable story draft: we introduce a dynamic storyline modeling that transforms user prompts into shot sequences through a two-stage process with self-validation mechanisms that enforce narrative coherence by rejecting candidates violating cinematic principles.
Additionally, to resolve visual inconsistency, we introduce identity-aware cross-shot propagation that extracts multi-aspect portrait schemata from narrative elements, ranging from different avatars in the same story to different traits of the same identity following the development of the story. This system handles both inter-story avatar variations and intra-story identity evolution, generating identity-preserving portrait (IPP) tokens. These IPP tokens guide keyframe generation through hierarchical feature injection in pretrained diffusion models [45, 40, 47, 55], ensuring style uniformity and identity fidelity across shots.
We encode each keyframe into latent space and utilize a video diffusion model to refine a noise map into video latent codes, representing $k$ frames of shot video, conditioned on the keyframe latent and the corresponding textual latent. To address transition artifacts, we design a cross-shot transition mechanism with a FIFO-like [30] latent reset strategy that processes adjacent shots’ features at the boundary, ensuring seamless transitions and maintaining visual continuity across shots while preserving the logical coherence established in the storyline preparation.
Additionally, current evaluation protocols for multi-shot video generation remain inadequate due to the absence of dedicated datasets and task-specific metrics. To address this limitation, we propose four novel quantitative metrics for systematic assessment:
  • Within-Shot Face Consistency (WS-FC): facial similarity between frames in a single shot.
  • Cross-Shot Face Consistency (CS-FC): identity distance across shots.
  • Within-Shot Style Consistency (WS-SC): style similarity between frames in a single shot.
  • Cross-Shot Style Consistency (CS-SC): stylized bias across shots.
Experimental results demonstrate VGoT’s superiority over state-of-the-art (SOTA) methods across all metrics. Quantitative comparisons reveal 20.4% and 17.4% improvements in WS-FC and WS-SC, respectively, compared to previous SOTA baselines. For cross-shot metrics, VGoT achieves 2.9$\times$ higher in CS-FC and 106.6% higher in CS-SC over baselines. Human evaluations (Table 2) confirm these findings, with VGoT receiving 66.7% ”Good” ratings for cross-shot consistency versus 27.2% for competitors.
The principal contributions of this work can be summarized as follows:
  • Automated Multi-Shot Generation Framework: We present VGoT, the first end-to-end system that generates story-coherent multi-shot videos from single sentence inputs, obviously reducing manual intervention.
  • Three-Core Solution Architecture: We propose a structured four-module pipeline addressing
@@
      Narrative Fragmentation through LLM-powered story decomposition
@@
      Visual Inconsistency via identity-aware propagation
@@
      Transition Artifacts using adjacent latent transition mechanisms
  • Multi-Shot Evaluation Protocol: We design a new multi-shot video assessment protocol featuring hierarchical consistency measurement with four quantitative metrics covering face/style consistency across both within-shot and cross-shot.
@@ 2 Related Work
Video generation has made significant strides following the great success of diffusion models, leading to two important research categories: long video synthesis and multi-shot story generation. These areas focus on generating high-quality, consistent videos either as extended single shots or as coherent sequences across multiple scenes.
Long Video Synthesis. Long video synthesis has advanced through diffusion-based methods and autoregressive approaches. Diffusion models based on Stable Diffusion [40] utilize iterative refinement to generate visually consistent frames and have been effective for short sequences [22, 5, 7, 8, 52, 50, 6, 21, 54, 59, 17, 29, 36, 13, 26, 44, 58, 61, 37, 35]. However, they struggle with maintaining coherence over extended video lengths. Autoregressive methods [15, 32, 56] predict frames sequentially but often face error accumulation, making long-term consistency challenging and computationally expensive. Additionally, training-free methods like FIFO-Diffusion [30] generate long sequences without training but lack mechanisms to manage transitions across shots, limiting their effectiveness in narrative-driven content. Overall, while these approaches achieve visual fidelity, they fail to ensure logical coherence across extended sequences. In contrast, VideoGen-of-Thought (VGoT) leverages a modular approach that includes cross-shot smoothing mechanisms to ensure both visual consistency and narrative coherence, offering a more holistic solution for generating long-form videos.
Multi-Shot Video Generation. Multiple Shot Story Generation focuses on maintaining narrative coherency across distinct scenes, and existing approaches face critical limitations in automation scalability. Animate-a-Story [23] uses retrieval-augmented generation to ensure visual consistency but struggles with maintaining logical narrative transitions. MovieDreamer [60] requires extensive manual scripting (character profiles, scene details, etc.) and image curation. DreamFactory [51] demands repetitive adjustment and special documents for each story in its multi-agent system across stories. Flexifilm [34] and StoryDiffusion [62] introduce conditional adaptability but still necessitate manual consistency fixes between shots. These methods share a common bottleneck: heavy reliance on human intervention for narrative and visual coherence. VGoT breaks this paradigm through systematic automation: 1) Narrative Fragmentation → LLM-driven story decomposition; 2) Visual Inconsistency → Story-derived identity propagation; 3) Transition Artifacts → Boundary-aware latent processing. This structured approach enables director-level storytelling from single prompts with 10× fewer manual interventions than prior works [60, 51].
@@ 3 Preliminaries and Problem Formulation
Figure 2: The FlowChart of VideoGen-of-Thought. Left: Shot descriptions are generated based on user prompts, describing various attributes such as character details, background, relations, and camera pose. Pre-shot descriptions provide a broader context for the upcoming scenes. Middle Top: Keyframes are generated using a text-to-image diffusion model conditioned with identity-preserving (IP) embeddings, which ensures consistent representation of characters throughout the shots. IP portraits help maintain visual identity consistency. Right: The shot-level video clips are generated from keyframes, followed by shot-by-shot transition inference to ensure temporal consistency across different shots. This collaborative framework ultimately produces a cohesive narrative-driven video.
@@ 3.1 Preliminaries
Diffusion models [25, 45, 46] are generative models trained to approximate data distributions $p(x)$ through iterative denoising of random noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. The forward process gradually adds noise according to a variance schedule $\beta_{t}$:
$q(x_{t}|x_{t-1})=\mathcal{N}\left(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I}\right)$ (1)
over $T$ timesteps, producing progressively noisier latents $\{x_{t}\}_{t=1}^{T}$. The reverse process learns a parameterized model $\mu_{\theta}$ to reconstruct $x_{0}$ through transitions:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}\left(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)\right)$ (2)
where $\mu_{\theta}$ and $\Sigma_{\theta}$ denote the predicted mean and variance. Training minimizes the noise prediction error via:
$\mathcal{L}_{\text{uncond}}=\mathbb{E}_{x_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I}),t}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right]$ (3)
Latent diffusion models [40] map this process to a compressed space using a VAE [31], enabling conditional generation through cross-attention:
@@
where $Q=W_{Q}\mathcal{E}(x_{t})$, $K=W_{K}\tau_{\theta}(y)$, and $V=W_{V}\tau_{\theta}(y)$ for text embeddings $y$. Video diffusion models [22, 5] process frame sequences $\{z^{f}\}_{f=0}^{F-1}\in\mathbb{R}^{h\times w\times d}$ through temporal-aware denoising networks $\epsilon_{\theta}(z_{t},t,c)$. FIFO-Diffusion [30] extends this through queue-based latent processing:
$Q_{k}=\{z^{f}\}_{f=f_{k}}^{f_{k}+n}\leftarrow\Phi(Q_{k},\tau_{k},c;\epsilon_{\theta})$ (5)
where $\Phi$ denotes the DDIM sampler [45]. While effective for frame continuity, this approach struggles with: (1) abrupt shot transitions that disrupt queue coherence, and (2) integration of image-based conditions. Our framework addresses these through reset mechanisms in adjacent latent-space processing.